Ben is 70 years old and Ishaan is 18 years old.
How many years will it take until Ben is only 3 times as old as Ishaan?

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

x = years in future when ben is three times ishaan's age
---
70 + x = 3(18 + x)
70 + x = 54 + 3x
2x = 16
x = 8
---
answer:
in 8 years ben will be three times ishaan's age
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

Ben is 70 years old and Ishaan is 18 years old.
How many years will it take until Ben is only 3 times as old as Ishaan?
------------------------
Let that number of years be "x"::
----------------------
Equation:
70+x = 3(18+x)
70+x = 54 + 3x
2x = 16
x = 8 years
----------------
Cheers,
Stan H.
=============